OpenStudy (anonymous):

The measures of two of the sides of an equilateral triangle are 3x + 15 in. and 7x – 5 in. What is the measure of the third side in inches?

jimthompson5910 (jim_thompson5910):

key thing here is that we have an equilateral triangle

jimthompson5910 (jim_thompson5910):

so all sides are equal

OpenStudy (anonymous):

has three equal sides. 3x + 15 = 7x - 5 4x = 20 x = 5 so the third side is: 3*5 + 15 = 30. <=====

jimthompson5910 (jim_thompson5910):

so 3x+15 = 7x-5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh, well thanks guys :)